# Model Breadcrumbs: Scaling Multi-Task Model Merging with Sparse Masks

### v1

@@ Abstract
The rapid development of AI systems has been greatly influenced by the emergence of foundation models. A common approach for targeted problems involves fine-tuning these pre-trained foundation models for specific target tasks, resulting in a rapid spread of models fine-tuned across a diverse array of tasks. This work focuses on the problem of merging multiple fine-tunings of the same foundation model derived from a spectrum of auxiliary tasks. We introduce a new simple method, Model Breadcrumbs, which consists of a sparsely defined set of weights that carve out a trajectory within the weight space of a pre-trained model, enhancing task performance when traversed. These breadcrumbs are constructed by subtracting the weights from a pre-trained model before and after fine-tuning, followed by a sparsification process that eliminates weight outliers and negligible perturbations. Our experiments demonstrate the effectiveness of Model Breadcrumbs to simultaneously improve performance across multiple tasks. This contribution aligns with the evolving paradigm of updatable machine learning, reminiscent of the collaborative principles underlying open-source software development, fostering a community-driven effort to reliably update machine learning models. Our method is shown to be more efficient and unlike previous proposals does not require hyperparameter tuning for each new task added. Through extensive experimentation involving various models, tasks, and modalities we establish that integrating Model Breadcrumbs offers a simple, efficient, and highly effective approach for constructing multi-task models and facilitating updates to foundation models. [^1] The code to reproduce our results is publicly available at: https://github.com/rezazzr/breadcrumbs
@@ 1 Introduction
In recent years, foundational models [3] have become instrumental tools, exhibiting unprecedented efficacy across multiple domains. These models are characterized by their extensive scale, generality, and capacity to learn and generalize knowledge from vast datasets, offering promising solutions to a diverse range of problems. The inherent ability of foundational models to be fine-tuned has led to advancements in natural language understanding (NLP) [41, 42, 13, 31, 45, 29], computer vision [43, 48, 33, 24, 5], and other related fields [49, 60, 50].
@@
To address the challenges of scalability, practical constraints, and unlock the untapped potential of the growing pool of publicly available fine-tuned models, recent developments in neural network weight averaging techniques have gained attention [22, 38, 47, 20, 57, 56, 6, 15]. These approaches enable the practitioners to re-purpose the increasingly valuable publicly available fine-tuned models.
Closer to our approach, Task Vectors were introduced by Ilharco et al. [20]. In their method, a foundation model is refined by incorporating the scaled average of the differences between multiple fine-tuned models and the foundation model. This allows for the creation of a multi-task model without the need for additional training or access to the original training data. However, despite its potential, the Task Vector method encounters limitations when dealing with numerous tasks. This is mainly due to its dependence on hyperparameter tuning through validation set performance, a process that becomes computationally impractical at scale, coupled with an increasing accumulation of noise as more tasks are merged to the foundation model.
@@ 2 Related Work
Model Merging
Recent studies in the literature have explored the merging of models trained from scratch with different initializations [1, 52]. One of the main challenges in this type of model merging is aligning the models before the actual merger. Therefore, research in this branch primarily focuses on finding permutations between networks to bring them into alignment with a reference model, enabling the subsequent merger of the two models in weight space. Moreover, this branch typically focuses on the merging of only two models. Our work, on the other hand, distinguishes itself from this line of research, as we concentrate on the model merging of networks that share the same initialization, specifically initialized by a foundation model. Furthermore, our investigation extends beyond the merging of just two models, exploring the dynamics when multiple models are involved in the merger process.
@@
Wortsman et al. [56] demonstrated that models fine-tuned on the same dataset with different hyperparameters can be combined together in a weighted average to yield an overall higher performing model. Unlike our work they did not consider merging models from different datasets and tasks. Choshen et al. [6] merges models from multiple trained models in order to create a better pretrained model for downstream tasks. Unlike our work they do not demonstrate or study the creation of multi-task ability through the merging. Matena and Raffel [35] considered merging of multiple fine-tuned models originating from the same pre-trained model, trained on diverse datasets. The merger operation combines a series of fine-tuned models using a weighted average determined by the Fisher information matrix [36]. However, computing the Fisher information matrix, as well as finding other required hyperparameters for this approach, becomes increasingly computationally expensive as the number of models to be merged grows. Therefore, it faces challenges when applied at scale. In contrast, our approach is computationally efficient, and as we will show in Section 4, its hyperparameters exhibit the ability to generalize to the scanrio where a large number of models are to be merged. A closely related work to ours is the study by Ilharco et al. [20], in which they introduced Task Vectors. Their method involves a scale averaging of the models fine-tuned for various tasks, originated from the same pre-trained model, to generate multi-task models. However, their approach necessitates a validation set for each new task, which adds complexity and computational overhead, coupled with an increasing accumulation of noise as more tasks are merged to the foundation model.
Federated Learning
The concept of initiating learning with a pre-trained model has been explored in the federated learning literature, as seen in recent works such as [39, 28]. These studies focused on a single downstream task where data is distributed across multiple clients. In their approach, each client periodically aggregates models during the training process. It’s important to note that this differs from our approach, which deals with multi-task learning involving multiple downstream tasks rather than a single task distributed across clients.
@@ 3 Model Breadcrumbs Framework
Figure 1: Method overview. We start with a foundational model that has undergone fine-tuning on various tasks. Next, we build a fine-tuning trajectory for each fine-tuned model by subtracting the pre-trained model weights from each of the fine-tuned models. We then, at each layer, apply a masking operation over the absolute value of the the resulting trajectory, eliminating both outliers and small values. Finally, these masked trajectories are aggregated and combined with the reference pre-trained model to create a unified multi-task model.
@@
To start generating Model Breadcrumbs, we begin with a pre-trained foundation model that has undergone fine-tuning for various auxiliary tasks. Denoting the weights of the foundation model as $\theta$, after fine-tuning on a specific task $t$, the weights are transformed into $\theta^{\prime}_{t}$. The initial step involves creating task direction by calculating the weight differences between $\theta^{\prime}_{t}$ and $\theta$, resulting in $\theta^{d}_{t}$.
@@
Note that $\theta^{d}_{t}$ contains both (a) large outliers, indicating substantial deviations from the pre-trained starting point, and (b) negligible differences representing minor perturbations from the foundation model’s weights. The presence of these extremes can impact the effectiveness of the resulting multi-task model upon merging. To address this concern, we implement a sparsification process that filters out both large outliers and small perturbations.
We mask out the extreme tails of the absolute magnitude distribution of $\theta^{d}_{t}$, using $\gamma$ and $\beta$ as thresholds for the right and left tails, respectively. Thus, a weight $w_{i}$ within layer $L$ (of size $n$) of $|\theta^{d}_{t}|$ undergoes masking if it meets either of the following conditions, where $L=\left\{w_{1},w_{2},\ldots,w_{n}\right\}$ and its elements are ordered as $w_{1}\leq w_{2}\leq\ldots\leq w_{n}$.
$\begin{cases}\text{if}\quad w\in\left\{w_{\gamma},w_{\gamma+1},\ldots,w_{n}\right\}\\ \text{if}\quad w\in\left\{w_{1},w_{2},\ldots,w_{\beta}\right\}\\ \end{cases}$ (2)
The masked weights are set to zero in $\theta^{d}_{t}$ or returned to their respective pre-training weights in $\theta^{\prime}_{t}$. We denote the resulting mask as $m^{\beta,\gamma}_{t}$. The remaining weights in $\theta^{d}_{t}$ remain unmasked.
Next, we apply the mask $m^{\beta,\gamma}_{t}$ to the task direction $\theta^{d}_{t}$. We now have a set of weight differences that define a trajectory within the weight space of the foundation model. Traversing this trajectory allows us to effectively transfer the knowledge accumulated during fine-tuning across tasks, while filtering out the harmful perturbations. For a total of $T$ tasks, we assemble a multi-task model $\theta^{*}$ by following the trajectories defined by the Model Breadcrumbs with a specific strength parameter $\alpha$. The formation of this multi-task model is expressed in Eq. 3. Algorithm 1 describes the overall procedure of the Model Breadcrumbs merging strategy.
@@
       $\theta^{d}_{t}\leftarrow\theta^{\prime}_{t}-\theta$ for $\mathrm{layer}\in\mathrm{Layers}(\theta)$  do
@@
             $p\leftarrow|\theta^{d}_{t,\mathrm{layer}}|$ /* Generate a mask for top k percent of the weights */
             $m_{t,\mathrm{layer}}^{\gamma}\leftarrow\mathrm{mask\_topk\_percent}(p,k=\gamma)$ /* Generate a mask for the bottom k percent of the weights */
             $m_{t,\mathrm{layer}}^{\beta}\leftarrow\mathrm{mask\_bottomk\_percent}(p,k=\beta)$ $m^{\beta,\gamma}_{t,\mathrm{layer}}\leftarrow\mathrm{merge\_masks}(m_{t,\mathrm{layer}}^{\beta},m_{t,\mathrm{layer}}^{\gamma})$
@@
$\theta^{*}\leftarrow\theta+\alpha\sum_{t\in T}m^{\beta,\gamma}_{t}.\theta^{d}_{t}$ return $\theta^{*}$
@@ 4 Experiments
In this section, we conduct a series of experiments to evaluate the Model Breadcrumbs framework. Our experiments focus on the following key aspects: 1. Merging Model Breadcrumbs: We incrementally add tasks, totalling 8 in our investigation, to assess the scalability and performance of merged Model Breadcrumbs as the number of tasks increases. 2. Generalization of Hyperparameters: We explore how the hyperparameters introduced by Model Breadcrumbs—$\alpha$, $\beta$, and $\gamma$—generalize over the number of datasets. 3. Effect of Scale: We investigate the impact of the scale and complexity of the foundation models on the Model Breadcrumbs’ adaptability and robustness. 4. Enhancement of Fine-tuned Models: We examine the potential of enhancing the performance a fine-tuned model by merging related tasks into it. 5. Ablation Study: We study the importance of the design choices introduced by Model Breadcrumbs for successful and competitive model merging.
@@ 4.1 Data, Metrics, and Models
In our analysis, we follow the benchmarks and settings used by [20] for a more meaningful comparison with existing reports. We present results using normalized accuracy, a metric calculated as the ratio of accuracy achieved by the merged model to that of the fine-tuned model.
@@
It is noteworthy that the fine-tuned model establishes the upper bound with a normalized accuracy value of 1. Subsequently, the concept of average normalized accuracy is introduced, representing the mean normalized accuracy across multiple tasks. In Section 4.2, 4.3, 4.4, and 4.6, we assess our findings using an extensive set of 8 datasets: Cars [25], DTD [7], EuroSAT [18], GTSRB [19], MNIST [27], RESISC45 [4], SUN397 [58], and SVHN [37]. We fine-tune various CLIP models [43] (ViT-B-32, ViT-B-16, and ViT-L-14) to explore model merging. For more information on the datasets and the fine-tuning process see Appendix A.
In Section 4.5, we apply our approach to the NLP domain, specifically investigating four GLUE tasks [53] (MRPC [14], RTE [53], CoLA [54], and SST-2 [51]) based on the benchmarks used by [20, 56]. Our process involves fine-tuning the T5-base model [44] on these datasets and subsequently merging publicly available fine-tuned models from other datasets (IMDB [34], RACE [26], QASC [23], MultiNews [16], SQuAD [46], and CommonGen [30]) into each of them. Appendix B provides additional details on the datasets, the fine-tuning process, and the publicly available fine-tuned models we used in our experiments.
@@ 4.2 Merging Model Breadcrumbs
Figure 2: The solid line is the averaged normalized accuracy across all evaluation points. Each data point corresponds to an experiment involving a subset of the 8 tasks under study. Notably, it is evident that the Model Breadcrumbs (with 85% sparsity), consistently outperform the Task Vectors [20]. Specifically, in the experiment involving all eight tasks, the Model Breadcrumbs outperform the Task Vectors by a substantial margin of 8.33%.
@@
Figure 3: Validation Free Setting. For the ViT-B-32 model, we tune the hyperparameters of each method (breadcrumbs and task vectors) based on the first 1,2, or 3 tasks and add additional tasks using those hyperparameters (validation set free). Moreover, for the ViT-L-14 model, we only tune the hyperparameters for the 1 task scenario and evaluate on the additional tasks using those hyperparameters. We observe that breadcrumbs substantially outperforms task vectors in this setting.
@@
Breadcrumbs 82.35
Task Vectors [20] 77.37
@@
Random Sparsed Task Vectors 74.00
Table 1: The evaluation of the above merging strategies over 8 tasks using ViT-B-32 reveals the advantage of Breadcrumbs over other merging methods. Note that not only Fisher Merging [35] lags behind both Task Vectors [20] and Model Breadcrumbs, it also requires significantly more computational resources.
In this section, we explore the scalability and performance of merged Model Breadcrumbs as we progressively include tasks, reaching a total of 8 in our investigation, as detailed in Section 4.1. Merging enables the creation of multi-task models that can excel across various tasks simultaneously. This versatility is valuable both in scenarios where we have multiple privately fine-tuned models as well as in cases where we have access to publicly available fine-tuned models. This allows the extraction of existing knowledge from these models without the need for extra training or access to additional training data. Table 1 presents a comparison between Model Breadcrumbs with 85% sparsity, the recently proposed Task Vectors [20], and Fisher Merging [35] across 8 tasks, using ViT-B-32 model. Model Breadcrumbs outperforms all considered methods by a substantial margin. Fisher Merging [35] lags behind both Task Vectors [20] and Model Breadcrumbs, and notably, it requires significantly more computational resources. Therefore, we proceed with the rest of our studies without evaluating Fisher Merging.
In Figure 2, we assess all possible task subsets of the 8 tasks detailed in Section 4.1, amounting to a total of $256=2^{8}$ combinations, under two settings: 1. evaluation over all 8 tasks and, 2. evaluation only on the subset of tasks that have been observed. As we can see in Figure 1(a) merging Model Breadcrumbs (85% sparsity) results in superior multi-task models compared to the Task Vectors [20]. Furthermore, the performance gap between these two approaches increases as more tasks are observed, resulting in vastly superior multi-task models when more Model Breadcrumbs are available.
In Figure 1(b) we can see that for small task numbers the resulting merged model performs closely to that of the multiple fine-tuned models although the gap increases as more tasks are added. Model Breadcrumbs again prove to be more performance that Task Vectors [20] in this setting.

### v2

@@ Abstract
The rapid development of AI systems has been greatly influenced by the emergence of foundation models. A common approach for targeted problems involves fine-tuning these pre-trained foundation models for specific target tasks, resulting in a rapid spread of models fine-tuned across a diverse array of tasks. This work focuses on the problem of merging multiple fine-tunings of the same foundation model derived from a spectrum of auxiliary tasks. We introduce a new simple method, Model Breadcrumbs, which consists of a sparsely defined weight set that guides model adaptation within the weight space of a pre-trained model. These breadcrumbs are constructed by subtracting the weights from a pre-trained model before and after fine-tuning, followed by a sparsification process that eliminates weight outliers and negligible perturbations. Our experiments demonstrate the effectiveness of Model Breadcrumbs to simultaneously improve performance across multiple tasks. This contribution aligns with the evolving paradigm of updatable machine learning, reminiscent of the collaborative principles underlying open-source software development, fostering a community-driven effort to reliably update machine learning models. Our method is shown to be more efficient and unlike previous proposals does not require hyperparameter tuning for each new task added. Through extensive experimentation involving various models, tasks, and modalities we establish that integrating Model Breadcrumbs offers a simple, efficient, and highly effective approach for constructing multi-task models and facilitating updates to foundation models.  [^1] The code to reproduce our results is publicly available at: https://github.com/rezazzr/breadcrumbs
Keywords:
Model Merging Transfer Learning Foundation Models
@@ 1 Introduction
In recent years, foundational models [3] have become instrumental tools, exhibiting unprecedented efficacy across multiple domains. These models are characterized by their extensive scale, generality, and capacity to learn and generalize knowledge from vast datasets, offering promising solutions to a diverse range of problems. The inherent ability of foundational models to be fine-tuned has led to advancements in natural language processing (NLP) [43, 44, 14, 32, 45, 30], computer vision [42, 49, 34, 25, 5], and other related fields [50, 62, 51].
@@
To address the challenges of scalability, practical constraints, and unlock the untapped potential of the growing pool of publicly available fine-tuned models, recent developments in neural network weight averaging techniques have gained attention [23, 11, 39, 48, 21, 58, 57, 6, 16, 60]. These approaches enable the practitioners to re-purpose the increasingly valuable publicly available fine-tuned models.
Closer to our approach, Task Arithmetic were introduced by Ilharco et al. [21]. In their method, a foundation model is refined by incorporating the scaled average of the differences between multiple fine-tuned models and the foundation model. This allows for the creation of a multi-task model without the need for additional training or access to the original training data. However, despite its potential, the Task Arithmetic method [21] encounters limitations when dealing with numerous tasks. This is mainly due to its dependence on hyperparameter tuning through validation set performance, a process that becomes computationally impractical at scale, coupled with an increasing accumulation of noise as more tasks are merged to the foundation model.
@@ 2 Related Work
Model Merging Recent studies in the literature have explored the merging of models trained from scratch with different initializations [1, 53]. One of the main challenges in this type of model merging is aligning the models before the actual merger. Therefore, research in this branch primarily focuses on finding permutations between networks to bring them into alignment with a reference model, enabling the subsequent merger of the two models in weight space. Our work, on the other hand, distinguishes itself from this line of research, as we concentrate on the model merging of networks that share the same initialization, specifically initialized by a foundation model. Furthermore, our investigation is focused on the scalability of merging methods, exploring the dynamics when multiple models are involved in the merger process.
@@
Wortsman et al. [57] demonstrated that models fine-tuned on the same dataset with different hyperparameters can be combined together using a weighted average to yield an overall higher performing model. Unlike our work they did not consider merging models from different datasets and tasks. Choshen et al. [6] merged models from multiple trained models in order to create a better pretrained model to be used for downstream tasks. Unlike our work they do not demonstrate or study the creation of multi-task ability through the merging. Matena and Raffel [36] considered merging of multiple fine-tuned models originating from the same pre-trained model, trained on diverse datasets. The merger operation combines a series of fine-tuned models using a weighted average determined by the Fisher information matrix [37]. However, computing the Fisher information matrix, as well as finding other required hyperparameters for this approach, becomes increasingly computationally expensive as the number of models to be merged grows. Therefore, it faces challenges when applied at scale. In contrast, our approach is computationally efficient, and as we will show in Section 4, its hyperparameters exhibit the ability to generalize to the scenarios where a large number of models are to be merged.
A related study to ours is conducted by Ilharco et al. [21], introducing a method named Task Arithmetic for model merging. Their approach begins by forming Task Vectors, representing the weight differences between pre-trained and fine-tuned weights for each task. The merged model’s weights are then obtained by adding a scaled sum of these task vectors to the pre-trained weights. However, their approach necessitates a validation set for each new task, which adds complexity and computational overhead, coupled with an increasing accumulation of noise as more tasks are merged to the foundation model.
A concurrent study by Yadav et al. [60] presents a method named TIES. Like the Task Arithmetic method [21], TIES initially constructs a set of Task Vectors. These vectors undergo a masking process to eliminate interfering weights, identified as a percentage of overall weights with low magnitudes. The remaining unmasked weights undergo a sign alignment operation to determine their polarity. Finally, a scaled sum merges the task vectors with the pre-trained model. Our approach differs from TIES in two key aspects. Firstly, we apply masking to both very large and small magnitude weights of the task vectors to minimize interference, whereas TIES focuses solely on small magnitude weights. Secondly, our masking strategy employs layer-wise masking as opposed to overall masking. Notably, in the context of task vectors, overall masking of small magnitude weights typically targets weights in the early layers [36].
Federated Learning The concept of initiating learning with a pre-trained model has been explored in the federated learning literature, as seen in recent works such as [40, 29]. These studies focused on a single downstream task where data is distributed across multiple clients. In their approach, each client periodically aggregates models during the training process. It’s important to note that this differs from our approach, which deals with multi-task learning involving multiple downstream tasks rather than a single task distributed across clients.
@@ 3 Model Breadcrumbs Framework
Figure 1: Method overview. We start with a foundational model that has undergone fine-tuning on various tasks. Next, we build a fine-tuning trajectory for each fine-tuned model by subtracting the pre-trained model weights from each of the fine-tuned models (task vectors). We then, at each layer, apply a masking operation over the absolute value of the the resulting trajectory, eliminating both outliers and small values. Finally, these masked task vectors are aggregated and combined with the reference pre-trained model to create a unified multi-task model.
@@
To start generating Model Breadcrumbs, we begin with a pre-trained foundation model that has undergone fine-tuning for various auxiliary tasks. Denoting the weights of the foundation model as $\theta$, after fine-tuning on a specific task $t$, the weights are transformed into $\theta^{\prime}_{t}$. The initial step involves creating task vectors by calculating the weight differences between $\theta^{\prime}_{t}$ and $\theta$, resulting in $\theta^{d}_{t}$.
@@
Note that $\theta^{d}_{t}$ contains both (a) large outliers, indicating substantial deviations from the pre-trained starting point, and (b) negligible differences representing minor perturbations from the foundation model’s weights. The presence of these extremes can impact the effectiveness of the resulting multi-task model upon merging. To address this concern, we implement a masking process that filters out both large outliers and small perturbations.
In each layer $L$, we mask out the extreme tails of the absolute magnitude distribution of $\theta^{d}_{t}$, using $\gamma$ and $\beta$ as thresholds for the right and left tails, respectively. Let $w^{L}_{i}$ represent the index of the weights sorted by their absolute magnitude in layer $L$ and $i$ the order in the sort (lowest to highest). The mask $m_{L}^{\beta,\gamma}$ for the layer $L$ is defined as:
$m_{L}^{\beta,\gamma}[w_{i}]=\begin{cases}0&\text{if }i\leq\beta\text{ or }i\geq\gamma\\ 1&\text{otherwise}\end{cases}$ (2)
The masked weights are set to zero in $\theta^{d}_{t}$ or returned to their respective pre-training weights in $\theta^{\prime}_{t}$. Aggregating $m_{L}^{\beta,\gamma}$ over all layers, for task $t$, results in the final mask $m^{\beta,\gamma}_{t}$.
Next, we apply the mask $m^{\beta,\gamma}_{t}$ to the task vectors $\theta^{d}_{t}$. We now have a set of weight differences that define a trajectory within the weight space of the foundation model. Traversing this trajectory allows us to effectively transfer the knowledge accumulated during fine-tuning across tasks, while filtering out the harmful perturbations. For a total of $T$ tasks, we assemble a multi-task model $\theta^{*}$ by following the trajectories defined by the Model Breadcrumbs with a specific strength parameter $\alpha$. The formation of this multi-task model is expressed in Eq. 3. Algorithm 1 describes the overall procedure of the Model Breadcrumbs merging strategy.
@@
      $\theta^{d}_{t}\leftarrow\theta^{\prime}_{t}-\theta$
      for $\mathrm{layer}\in\mathrm{Layers}(\theta)$  do
@@
            $p\leftarrow|\theta^{d}_{t,\mathrm{layer}}|$
             /* Generate a mask for top k percent of the weights */
            $m_{t,\mathrm{layer}}^{\gamma}\leftarrow\mathrm{mask\_topk\_percent}(p,k=\gamma)$
             /* Generate a mask for the bottom k percent of the weights */
            $m_{t,\mathrm{layer}}^{\beta}\leftarrow\mathrm{mask\_bottomk\_percent}(p,k=\beta)$
            $m^{\beta,\gamma}_{t,\mathrm{layer}}\leftarrow\mathrm{merge\_masks}(m_{t,\mathrm{layer}}^{\beta},m_{t,\mathrm{layer}}^{\gamma})$
@@
$\theta^{*}\leftarrow\theta+\alpha\sum_{t\in T}m^{\beta,\gamma}_{t}.\theta^{d}_{t}$
return $\theta^{*}$
@@ 4 Experiments
In this section, we conduct a series of experiments to evaluate the Model Breadcrumbs framework. Our experiments focus on the following key aspects: 1. Merging Model Breadcrumbs: We incrementally add tasks, totalling 8 in our investigation, to assess the scalability and performance of merged Model Breadcrumbs as the number of tasks increases. 2. Generalization of Hyperparameters: We explore how the hyperparameters introduced by Model Breadcrumbs—$\alpha$, $\beta$, and $\gamma$—generalize over the number of datasets. 3. Effect of Scale: We investigate the impact of the scale and complexity of the foundation models on the Model Breadcrumbs’ adaptability and robustness. 4. Target Task Improvement: We examine the potential of enhancing the performance a fine-tuned model on a target task by merging related tasks into it. 5. Ablation Study: We study the importance of the design choices introduced by Model Breadcrumbs for successful and competitive model merging.
@@ 4.1 Data, Metrics, and Models
In our analysis, we follow the benchmarks and settings used by Ilharco et al. [21] for a more meaningful comparison with existing reports. We present results using normalized accuracy, a metric calculated as the ratio of accuracy achieved by the merged model to that of the fine-tuned model.
@@
It is noteworthy that the fine-tuned model establishes the upper bound with a normalized accuracy value of 1. Subsequently, the concept of average normalized accuracy is introduced, representing the mean normalized accuracy across multiple tasks. In Section 4.2, 4.3, 4.4, and 4.6, we assess our findings using an extensive set of 8 datasets: Cars [26], DTD [7], EuroSAT [19], GTSRB [20], MNIST [28], RESISC45 [4], SUN397 [59], and SVHN [38]. We fine-tune various CLIP models [42] (ViT-B-32, ViT-B-16, and ViT-L-14) to explore model merging. For more information on the datasets and the fine-tuning process see Appendix 0.A.
In Section 4.5, we apply our approach to the NLP domain, specifically investigating four GLUE tasks [54] (MRPC [15], RTE [54], CoLA [55], and SST-2 [52]) based on the benchmarks used by [21, 57]. Our process involves fine-tuning the T5-base model [46] on these datasets and subsequently merging publicly available fine-tuned models from other datasets (IMDB [35], RACE [27], QASC [24], MultiNews [17], SQuAD [47], and CommonGen [31]) into each of them. Appendix 0.B provides additional details on the datasets, the fine-tuning process, and the publicly available fine-tuned models we used in our experiments.
@@ 4.2 Merging Model Breadcrumbs
Figure 2: The solid line is the averaged normalized accuracy across all evaluation points. Each data point corresponds to an experiment involving a subset of the 8 tasks under study. Notably, it is evident that the Model Breadcrumbs (with 90% sparsity), consistently outperform the Task Arithmetic [21]. Specifically, in the experiment involving all eight tasks, the Model Breadcrumbs outperform the Task Arithmetic by a substantial margin of 5.7%.
@@
Figure 3: Validation Free Setting. For the ViT-B-32 model, we tune the hyperparameters of each method (Breadcrumbs and Task Arithmetic) based on the first 1, 2, or 3 tasks and add additional tasks using those hyperparameters (validation set free). For the ViT-L-14 model, the Breadcrumbs method was only tune for the 1 task scenario and evaluate on the additional tasks using those hyperparameters, though the Task Arithmetic appraoch was given more chances to adjust its hyperparameters (task 1, 2, and 3). We observe that Breadcrumbs substantially outperforms task vectors in this setting.
Table 1: The evaluation of the above merging strategies over 8 tasks using ViT-B-32 reveals the advantage of Breadcrumbs over other merging methods. Note that not only Fisher Merging [36] lags behind both Task Arithmetic [21] and Model Breadcrumbs, it also requires significantly more computational resources.
@@
Breadcrumbs 83.35
Task Arithmetic [21] 77.66
@@
Random Sparsed Task Arithmetic 74.00
In this section, we explore the scalability and performance of merged Model Breadcrumbs as we progressively include tasks, reaching a total of 8 in our investigation, as detailed in Section 4.1. Merging enables the creation of multi-task models that can excel across various tasks simultaneously. This versatility is valuable both in scenarios where we have multiple privately fine-tuned models as well as in cases where we have access to publicly available fine-tuned models. This allows the extraction of existing knowledge from these models without the need for extra training or access to additional training data. Table 1 presents a comparison between Model Breadcrumbs with 90% sparsity ($\beta=90\%$, $\gamma=99\%$), the recently proposed Task Arithmetic [21], and Fisher Merging [36] across 8 tasks, using ViT-B-32 model. Model Breadcrumbs outperforms all considered methods by a substantial margin. Fisher Merging [36] lags behind both Task Arithmetic [21] and Model Breadcrumbs, and notably, it requires significantly more computational resources. Therefore, we proceed with the rest of our studies without evaluating Fisher Merging.
In Figure 2, we assess all possible task subsets of the 8 tasks detailed in Section 4.1, amounting to a total of $256=2^{8}$ combinations, under two settings: 1. evaluation over all 8 tasks and, 2. evaluation only on the subset of tasks that have been observed. As we can see in Figure 2(a) merging Model Breadcrumbs (90% sparsity) results in superior multi-task models compared to the Task Arithmetic [21]. Furthermore, the performance gap between these two approaches increases as more tasks are observed, resulting in vastly superior multi-task models when more Model Breadcrumbs are available.
In Figure 2(b) we can see that for small task numbers the resulting merged model performs closely to that of the multiple fine-tuned models although the gap increases as more tasks are added. Model Breadcrumbs again prove to be more performance that Task Arithmetic [21] in this setting.